-(5y+5)-(-4y-7)=-5
Please help me solve this equation

Answer by HyperBrain(694)

-(5y+5)-(-4y-7)=-5
-5y+4y-5+7=-5
-y+2=-5
-y=-7
y=7
Power up,
HyperBrain!
Stop this mess! Click here